When the Hendersons went to dinner, their total bill, including tax, was $43.74. If the sales tax rate is 8%, what was the original price of the dinner?
Answer by kcanil08(11)

let their total bill was x.
then
x + 8% of x = 43.74
x + 8/100*x = 43.74
100x + 8x =43.74*100
x= 4374/108
x= 40.5